OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

OpenStudy (amistre64):

use the dot product ....\[cos~\alpha=\frac{\vec u \cdot \vec v}{|\vec u||\vec v|}\]

OpenStudy (amistre64):

inverse cosine to account for the angle ....

OpenStudy (anonymous):

so do I do something like subtract the x(u) from the x(v) to get the u ontop?

OpenStudy (amistre64):

i just stack u and v, multiply down, and add across u = <-5, -4> v = <-4, -3> --------- 20 + 12 = u.v

OpenStudy (amistre64):

now divide by the lengths ...

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

the lengths of what? I'm sorry

OpenStudy (amistre64):

the lengths of the vectors of course ....

OpenStudy (amistre64):

how long is u? how long is v ?

OpenStudy (amistre64):

by the pythag thrmt:\[|\vec a|=\sqrt{(\vec a \cdot \vec a)}\]

OpenStudy (anonymous):

what would the a stand for in this instance

OpenStudy (amistre64):

some generic vector, any vector. It is a property that is shared by all vectors and so the vectors in this problem use it as well. |v| v = <-4, -3> v = <-4, -3> -------------- sqrt( 16 + 9 )

OpenStudy (anonymous):

okay now we just do it for the other, then multiply it

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

5 * sqrt(41)

OpenStudy (anonymous):

?

OpenStudy (amistre64):

25 16 ---- 41 yep

OpenStudy (amistre64):

so, we have:\[cos~\alpha=\frac{20+12}{5\sqrt{41}}\] solve for alpha

OpenStudy (anonymous):

= .99951

OpenStudy (anonymous):

?

OpenStudy (amistre64):

cos a = .99951 a = cos^-1 (.99951), but this gets us radians multiply by pi/180 to convert to degrees

OpenStudy (amistre64):

i spose it depends on if your calculator is in degree mode or radian mode ... degree mode gives out 1.7937...

OpenStudy (anonymous):

okay, thanks!

OpenStudy (amistre64):

youre welcome